In the reaction c4h10 + o2= co2 + h2o 172.5 g of c4h10 reacts with excess o2. How many grams of co2 will be produced?
1
Expert's answer
2018-12-07T01:51:19-0500

Answer on Question 83627, Chemistry/ Organic Chemistry

In the reaction c4h10 + o2= co2 + h2o 172.5 g of c4h10 reacts with excess o2. How many grams of co2 will be produced?

Solution

2C4H10 +13 O2 = 8CO2 +10H2O

Find number of moles of C4H10:

n=m/M

M(C4H10) = 12.01*4 + 1.01*10 = 58.14 (g/mol)

n(C4H10)= 172.5 g/ 58.14 g/mol = 2.967 moles

According to equation 2 moles of C4H10 produce 8 moles of CO2

We have 2.967 moles of C4H10 that produce x moles of CO2

Solve the proportion:

2/2.967 = 8/x

x= 11.87

n(CO2) = 11.87 moles

m= M*n

M(CO2) = 12.01 + 16.00*2 = 44.01 (g/mol)

m(CO2) = 44.01 g/mol *11.87 moles = 522.4 g

Answer: 522.4 g